Use the observations made from the four experiments shown in the table below to determine the...

  1. Use the observations made from the four experiments shown in the table below to determine the rate law for the reaction 2 NO2(g) + F2(g)   2NO2F(g).  Find the exponents, x and y, and the rate constant, k, and its units (you only need to use Experiment 1 values to find k).    Rate = k [NO2]x [F2]y   (6 points)

______            ______            ___________________________      ________________

x                                  y                                  Rate Law                                                                           k, and its units

            

Experiment

Initial Rate (mole/L.s)

Initial [NO2] (mole/L)

Initial [F2] (mole/L)

1

0.0255

0.100

0.100

2

0.0510

0.200

0.100

3

0.102

0.200

0.200

4

0.408

0.400

0.400
Answer #1

Solution:-

The rate law for the reaction is:

Rate = k [NO2]x [F2]y

Substituting the values of rate, [NO2] and [F2] for various sets of experiments, we have

0.0255 = k [0.100]x [0.100]y ..................(i) [For Experiment 1]

0.0510 = k [0.200]x [0.100]y ..................(ii) [For experiment 2]

0.102 = k [0.200]x [0.200]y ..................(iii) [For experiment 3]

0.408 = k [0.400]x [0.400]y ...................(iv) [For experiment 4]

Dividing eq.(ii) by eq.(i), we get

(0.0510) / (0.0255) = [0.200]x / [0.100]x

2 = 2x

x = 1

Dividing eq.(iii) by eq.(ii), we get

(0.102) / (0.0510) = [0.200]y / [0.100]y

2 = 2y

y = 1

Therefore, the rate law is

Rate = k [NO2]1 [F2]1

Substituting the values of rate , [NO2 ] and [F2] for Experiment 1 , in the rate law, we get

0.0255 mol/ L-s = k [0.100 mol/L] [0.100 mol/L]

k = (0.0255 mol/L-s) / (0.100 mol/L × 0.100 mol/L)

(Rate constant) k = 2.55 mol-1 L s-1

------------------------------------------------------------------------------------